OpenStudy (anonymous):

solve each system by graphing and tell whether it is consistent or inconsistent if the system is consistent tell whether it is dependent or independent x-y=-4 x+3y=24

OpenStudy (turingtest):

I'm not going to solve it by graphing just sub one equation into the other\[x-y=-4\implies x=y-4\]\[ x+3y=24\]subbing the expression for x into the second equation we get\[y-4+3y=24\]solve for y if you can reduce the expression to 0=0 it has infinite solutions if you get a falsehood, like 1=2, then no solution if you get a value for y (which you can then use to find the value for x) that is the one unique solution

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

how do i get y

OpenStudy (turingtest):

how could you have possibly read all that I wrote in the 2 seconds this has been posted? I gave you a formula that contains only y by subbing in from the first equation to the second use that equation that I gave you to find y, it is a direct one-variable problem then use the first equation to find x

OpenStudy (anonymous):

i am a fast reader lol

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

the math i dont get

OpenStudy (turingtest):

you may have read it fast, but since you did not catch that I specifically told you how to find y, and even gave you the formula to find it, I felt perhaps you did not read carefully enough ;)

OpenStudy (anonymous):

I did read it i have to find the x

OpenStudy (anonymous):

x+3y=24

OpenStudy (turingtest):

do you have y yet?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

I am confused.. i give up..

OpenStudy (turingtest):

the first equation is\[x-y=-4\]that means that\[x=y-4\] agreed?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got to go in like 5 minutes

OpenStudy (turingtest):

the second equation is\[x+3y=24\] plug in the expression for x we agreed on above into this equation and you get\[(y-4)+3y=-4\]still with me?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

*typo\[(y-4)+3y=24\]I meant so you can solve this for y, so go ahead and do that

OpenStudy (anonymous):

4y

OpenStudy (turingtest):

=?

OpenStudy (anonymous):

=24

OpenStudy (turingtest):

no y-4+3y=24 4y=28 y=7

OpenStudy (turingtest):

now plug in y=7 to either of your other expression and you will get x

OpenStudy (anonymous):

y-4+3y=24?

OpenStudy (turingtest):

what about it? that is the formula we used to find y

OpenStudy (anonymous):

and put 7's where y is

OpenStudy (turingtest):

no, we need to go back to an equation that still has x in it so we can solve for x

OpenStudy (anonymous):

ok well i gotta go ..

OpenStudy (anonymous):

ill be back online in bout an hour

OpenStudy (turingtest):

x-y=-4 or x+3y=24 plug in y=7 to either and you get x later!